# Benchmarking Offline Goal-Conditioned RL

### v1

@@ Appendix C Offline GCRL Algorithms
Appendix D Implementation Details
@@
Table 7: Environment specifications.
@@
Table 8: Dataset specifications.
@@
D.1 Tasks and Datasets
D.2 Methods
We provide the full list of common hyperparameters in Table 9. We find that methods are more sensitive to policy extraction hyperparameters (e.g., the BC coefficient in DDPG+BC) (Park et al., 2024a), and report these in a separate table (Table 10). Specifically, for each method, we use the same value learning hyperparameters across the benchmark except for the discount factor $\gamma$ (Table 9), but individually tune the policy extraction hyperparameters (e.g., AWR $\alpha$ and DDPG+BC $\alpha$) for each dataset category (Tables 9 and 10).
@@ D.2 Methods
Appendix E Evaluation Goals and Per-Goal Benchmarking Results
Each task in OGBench provides five evaluation goals. We provide their full image descriptions in Figures 4, 5, 6, 7, 8, 9 and 10, and the full per-goal evaluation results in Tables 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22 and 23, which share the same format as Table 2.

### v2

@@ Appendix D Offline GCRL Algorithms
Appendix E Implementation Details
@@
E.1 Tasks and Datasets
E.2 Single-Task Variants
OGBench also supports standard (i.e., non-goal-conditioned) offline RL by providing single-task variants of locomotion and manipulation tasks. To convert a goal-conditioned task into a standard reward-maximizing task, we fix an evaluation goal and relabel the dataset with a semi-sparse reward function. This semi-sparse reward function is defined as the negative of the number of unaccomplished subtasks in the current state, and the episode immediately terminates when the agent completes all subtasks of the target evaluation goal. In locomotion environments, rewards are always $-1$ or 00, as there are no separate subtasks. In manipulation environments, rewards range between $-n_{\mathrm{task}}$ and 00, where $n_{\mathrm{task}}$ denotes the number of subtasks (e.g., in puzzle-4x6, $n_{\mathrm{task}}=24$ as there are $24$ buttons).
Each locomotion and manipulation task in OGBench provides five single-task variants that correspond to the five evaluation goals (Appendix F), resulting in a total of $410$ single-task tasks. They are named with the suffix “singletask-task[n]” (e.g., scene-play-singletask-task2-v0), where [n] denotes a number between $1$ and $5$ (inclusive). Among the five tasks in each environment, the most representative one is chosen as the “default” task, and is aliased by the suffix “singletask” without a task number. For example, in cube-double, the second task (standard double pick-and-place; see Figure 6) is set as the default task, and cube-double-play-singletask-v0 and cube-double-play-singletask-task2-v0 refer to the same task. Default tasks can be useful in various ways. For instance, one may report performance only on default tasks to reduce the computational burden, or may treat default tasks as a “training” task set for tuning hyperparameters while using the other four tasks as a “validation” task set. We provide the list of default tasks in Table 9.
While we do not provide a separate benchmarking result on the single-task environments, a benchmarking table of several representative offline RL algorithms on $50$ tasks can be found in the work by Park et al. (2025).
E.3 Oracle Representation Variants
OGBench also provides oracle representation variants of locomotion and manipulation tasks, denoted by the suffix “oraclerep” (e.g., antmaze-large-navigate-oraclerep-v0). These tasks provide low-dimensional oracle goal representations that contain only relevant information for fulfilling the goal success criterion. This corresponds to the $x$-$y$ coordinates of the agent (or the ball in antsoccer) in locomotion environments, and the positions of the cubes and the states of the objects in manipulation environments. The oraclerep tasks reduce the burden of goal representation learning, potentially helping diagnose the bottlenecks in goal-conditioned RL algorithms. We do not provide a separate benchmarking result for the oracle representation variants.
E.4 Methods
We provide the full list of common hyperparameters in Table 10. We find that methods are more sensitive to policy extraction hyperparameters (e.g., the BC coefficient in DDPG+BC) (Park et al., 2024a), and report these in a separate table (Table 11). Specifically, for each method, we use the same value learning hyperparameters across the benchmark except for the discount factor $\gamma$ (Table 10), but individually tune the policy extraction hyperparameters (e.g., AWR $\alpha$ and DDPG+BC $\alpha$) for each dataset category (Tables 10 and 11).
@@ E.4 Methods
Table 7: Environment specifications. See Table 8 for the dataset specifications. Note that the episode lengths of datasets and environments can be different.
@@
Table 8: Dataset specifications. See Table 7 for the environment specifications. Note that the episode lengths of datasets and environments can be different.
@@
Table 9: Designated default tasks for single-task environments. For single-task (singletask) variants, each environment provides five tasks corresponding to the five evaluation goals, with the most representative one chosen as the default task.
Environment Type Environment Default Task
pointmaze pointmaze-medium-v0 task1
pointmaze-large-v0 task1
pointmaze-giant-v0 task1
pointmaze-teleport-v0 task1
antmaze antmaze-medium-v0 task1
antmaze-large-v0 task1
antmaze-giant-v0 task1
antmaze-teleport-v0 task1
humanoidmaze humanoidmaze-medium-v0 task1
humanoidmaze-large-v0 task1
humanoidmaze-giant-v0 task1
antsoccer antsoccer-arena-v0 task4
antsoccer-medium-v0 task4
visual-antmaze visual-antmaze-medium-v0 task1
visual-antmaze-large-v0 task1
visual-antmaze-giant-v0 task1
visual-antmaze-teleport-v0 task1
visual-humanoidmaze visual-humanoidmaze-medium-v0 task1
visual-humanoidmaze-large-v0 task1
visual-humanoidmaze-giant-v0 task1
cube cube-single-v0 task2
cube-double-v0 task2
cube-triple-v0 task2
cube-quadruple-v0 task2
scene scene-v0 task2
puzzle puzzle-3x3-v0 task4
puzzle-4x4-v0 task4
puzzle-4x5-v0 task2
puzzle-4x6-v0 task2
visual-cube visual-cube-single-v0 task2
visual-cube-double-v0 task2
visual-cube-triple-v0 task2
visual-cube-quadruple-v0 task2
visual-scene visual-scene-v0 task2
visual-puzzle visual-puzzle-3x3-v0 task4
visual-puzzle-4x4-v0 task4
visual-puzzle-4x5-v0 task2
visual-puzzle-4x6-v0 task2
@@
Appendix F Evaluation Goals and Per-Goal Benchmarking Results
Each task in OGBench provides five evaluation goals. We provide their full image descriptions in Figures 4, 5, 6, 7, 8, 9 and 10, and the full per-goal evaluation results in Tables 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23 and 24, which share the same format as Table 2.